# On the origins of counterrotating stellar disks in TNG50. I

### v1

@@ 7 Summary and conclusions
In this work we have identified and analyzed the formation mechanisms of counterrotating stellar disks in galaxies selected from the Illustris TNG50 simulation. The sample focused only on central MW-mass galaxies, ranging in mass from $10^{11.5}$ to $10^{12.5}$M. We characterized these counterrotating components by selecting particles with circularities <0.7<-0.7 and located within the spatial extension of the stellar disc. Within the sample of 260 MW-mass galaxies, we identifed 26 cases with counterrotating stellar disk. Their CRDs contribute with, at least, 1 percent of the stellar mass of the galaxy. Our main results can be summarized as follows:
@@
  • Statistics: CRDs are rare. Only $\approx 10\%$ host a significant stellar counterrotating disk (CRD) when adopting a threshold of $M_{\rm CRD}/M_{\rm Disk}\geq 0.01$. CRDs represent a minor fraction of the overall stellar disk mass.
@@
  • Structure: CRDs are predominantly compact. Quantitatively, $\approx 88\%$ have $R^{\rm CRD}_{50}<0.25\,R_{\rm opt}$, and the distribution of sizes peaks at median $R^{\rm CRD}_{95}/R_{\rm opt}\sim 0.37$, although a minority are more extended; three systems display torus-like morphologies.
@@
  • Stellar origin: Nearly all CRDs are in-situ dominated. This definition includes stars formed inside the host from gas that may have been stripped from satellites. Ex-situ contributions are typically small ($\approx 73\%$ of CRDs have ex-situ <0.2<0.2; median ex-situ fraction $\approx 0.06$).
@@
  • Star-formation histories and triggers: CRDs are generally older than their co-rotating counterparts and exhibit bursty SFHs whose peaks frequently coincide with satellite pericenters or interactions. Thus, external perturbations often trigger star formation. These episodes may occur in retrograde or prograde orbits, which are subsequently transformed into a CRD through the later accretion of cold gas, giving rise to the present-day disk.
@@
  • Formation channels: Compact in-situ CRDs arise through multiple pathways: (i) early formation alongside bulge assembly; (ii) interaction-induced in-situ bursts; and (iii) rarer cases linked to smooth misaligned gas accretion. Extended in-situ CRDs, although fewer, can usually be traced to interactions and late accretion of misaligned gas that gives rise to a younger dominant corotating stellar disk. Ex-situ rich cases typically originate from one dominant (occasionally two) progenitor(s); these mergers often coincide with enhanced in-situ star formation in retrograde orbits.
@@
In a follow-up work we will characterize the mechanism that drives the late accretion of misaligned cold gas, giving rise to preferentially extended CRDs. We will also explore whether the identification of compact in-situ CRDs can be achieved thorugh the characterization of the stellar population abundances.
Acknowledgments
@@
MCB, FAG, AD and PBT acknowledges support from the ANID BASAL project FB210003. FAG acknowledges support from the ANID FONDECYT Regular grant 1251493. FAG, AD and PBT acknowledges support from the HORIZON MSCA-2021-SE-01 Research and Innovation Programme under the Marie Sklodowska-Curie grant agreement number 101086388. AD acknowledges support from the FONDECYT Postdoctorado No. 3250558. PBT acknowledges partial support from Fondecyt Regular ANID 1240465.
Appendix A Spatial distribution of components for galaxies hosting a significant CRD.
@@
Figures 10 and 11 show the face-on and edge-on projections of the stellar mass surface density for the 26 significant CRD galaxies.
@@
Figure 10: The face-on and edge-on projections of the stellar mass surface density for the 26 significant CRD galaxies. The galaxies are sorted as in figure 7. Here, we show all the stellar particles that belong to the host galaxy. The yellow solid and dotted-dashed circle represent the radii, $R_{95}^{\rm CRD}$ and $R_{50}^{\rm CRD}$, of the counterrotating disk, respectively, while the white dashed circle represents the optical radius, $R_{\rm opt}$, of the galaxy. In the top right corner of each panel the type is shown which are, Extended exsitu CRDs, Compact exsitu CRDs, Compact insitu CRDs, Extended insitu CRDs which are represented as yellow star, red rombuses, blue triangles, and orange circles respectively.
@@
Figure 11: As in 10 but for only counterrotating stellar disk particles, as described in Section 2.3

### v2

@@ 7 Summary and conclusions
In this work we have identified and analyzed the formation mechanisms of counterrotating stellar disks in galaxies selected from the Illustris TNG50 simulation. We limited our sample to only MW-mass galaxies, ranging in total mass from $10^{11.5}$ to $10^{12.5}$M. We characterized these counterrotating components by selecting particles with circularities of less than 0.7-0.7 and located within the spatial extension of the stellar disk. Within the sample of 260 MW-mass galaxies, we identified 26 cases with a counterrotating stellar disk. Their CRDs contribute to at least 1% of the stellar mass of the galaxy. Our main results are summarized as follows:
@@
  • Statistics: In the considered sample, CRDs are rare. Only $\approx 10\%$ of our sample hosts a significant CRD when adopting a threshold of $M_{\rm CRD}/M_{\rm Disk}\geq 0.01$. As expected in late-type galaxies, CRDs represent a minor fraction of the overall stellar disk mass.
@@
  • Structure: CRDs from our sample are predominantly compact. Quantitatively, $\approx 88\%$ have $R^{\rm CRD}_{50}<0.25\,R_{\rm opt}$, and the distribution of the sizes peaks at a median of $R^{\rm CRD}_{95}/R_{\rm opt}\sim 0.37$, although a minority are more extended. Notably, three systems display torus-like morphologies.
@@
  • Stellar origin: Nearly all the CRDs in our sample are in situ dominated. The definition we used includes stars formed inside the host from gas that may have been stripped from satellites. Ex situ contributions are typically small. Approximately 73% of the CRDs have ex situ fractions <0.2<0.2, and the median ex situ fraction is $\approx 0.06$.
@@
  • Star formation histories and triggers: CRDs from the analyzed sample are generally older than their corotating counterparts and exhibit bursty SFHs whose peaks frequently coincide with satellite pericenters or interactions. Thus, external perturbations often trigger star formation. These episodes may occur in retrograde or prograde orbits, which are subsequently transformed into a CRD through the later accretion of cold gas, which gives rise to the present-day disk.
@@
  • Formation channels: A variety of formation mechanisms appear in our sample of CRDs. Compact in situ CRDs arise through multiple pathways: (i) early formation alongside bulge assembly; (ii) interaction-induced in situ bursts; and (iii) rarer cases linked to smooth misaligned gas accretion. Extended in situ CRDs, although fewer, can usually be traced to interactions and late accretion of misaligned gas that gives rise to a younger dominant corotating stellar disk. ex situ rich cases typically originate from one dominant (occasionally two) progenitor(s); these mergers often coincide with enhanced in situ star formation in retrograde orbits.
@@
In a follow-up work we will characterize the mechanism that drives the late accretion of misaligned cold gas and preferentially gives rise to extended CRDs. We will also explore whether the identification of compact in situ CRDs can be achieved through the characterization of stellar population abundances.
Acknowledgements.
@@
We thank the anonymous referee for the insightful comments and constructive suggestions that helped improve this work. MCB, FAG, AD and PBT acknowledges support from the ANID BASAL project FB210003. FAG acknowledges support from the ANID FONDECYT Regular grant 1251493. FAG, AD and PBT acknowledges support from the HORIZON MSCA-2021-SE-01 Research and Innovation Programme under the Marie Sklodowska-Curie grant agreement number 101086388. AD acknowledges support from the FONDECYT Postdoctorado No. 3250558. PBT acknowledges partial support from Fondecyt Regular ANID 1240465. The authors acknowledge the use of ChatGPT to improve the clarity and readability of the manuscript.
Appendix A Spatial distribution of components for galaxies hosting a significant CRD
@@
Figures 12 and 13 show the face-on and edge-on projections of the stellar mass surface density for the 26 significant CRD galaxies. We show the scale size bar and the internal id on the top-left corner and left margin of each panel, respectively. The variety of configurations for the CRDs can also be appreciated.
@@
Figure 12: Face-on and edge-on projections of the stellar mass surface density for the 26 significant CRD galaxies. The galaxies are sorted as in figure 9. We show all the stellar particles that belong to the host galaxy. The solid and dotted-dashed yellow circles represent the radii, $R_{95}^{\rm CRD}$ and $R_{50}^{\rm CRD}$, of the coplanar counterrotating disk, respectively, while the white dashed circle represents the optical radius, $R_{\rm opt}$, of the galaxy. In the top-right corner of each panel, the type is shown, which are extended ex situ CRDs, compact ex situ CRDs, compact in situ CRDs, extended in situ CRDs. The types are respectively represented with a yellow star, red rhombuses, blue triangles, and orange circles.
@@
Figure 13: Same as in 12 but for only counterrotating stellar disk particles, as described in Section 2.3
Appendix B Orbital properties of infalling satellites in compact in situ CRDs.
Figure 14 shows the orbital properties for compact in situ CRDs that formed later, with their counterrotating stars born during one or more bursts triggered by satellite interactions (but without significant stellar accretion). The colors represent each orbit of a satellite and are the same as in Fig. 9. It can be appreciated how in each case at least one satellite initially had a misaligned orbit.
Figure 14: The orbit of the satellites that are in interaction with CR-11, CR-12, CR-18 and CR-20, respectively. Each column of panels is constrained to the lookback time window that has a peak in the SFH fo the CRD according to Fig. 9. Top panel: evolution of the satellite galactocentric distance from the host. Bottom panel: evolution of the angle between the angular momentum of the disk and the satellite’s orbital angular momentum. Orbits near $180^{\circ}$ are counterrotating, while those near $0^{\circ}$ are corotating.
Appendix C Differences with past studies on counterrotating disks
As mentioned in Sec. 1 of this paper, there have been a multitude of studies on the topic of counterrotating disks. Most of them, however, have focused on lenticular and elliptical galaxies where these types of structures are more common. Here, we compare a study by Khoperskov et al. (2021) with our work. In Figure 15, we show the distribution of the disk-to-total mass ratio (D/T) of the counterrotating galaxy sample selected in Khoperskov et al. (2021), according to Genel et al. (2015)’s catalog, and the sample in this work.
Figure 15: Distribution of D/T values of galaxy models with identified CRDs, extracted from the publicly available catalogs from the IllustrisTNG project Genel et al. (2015). The gray (dashed) and magenta (dash-dotted) histograms indicate the D/T distribution of CRDs galaxies identified in Khoperskov et al. 2021 and Bugueño et al. (2025), respectively.
We observed that the galaxies selected in this work are characterized by a higher D/T, meaning that they include late-type disk galaxies which are different from the early-type disk counterparts studies in Khoperskov et al. (2021). This also means that caution is required when comparing our results with observational studies, which have mainly studied CRDs in lenticular galaxies (Bevacqua et al. 2022; Gasymov et al. 2025; Katkov et al. 2013; Pillepich et al. 2018; Coccato et al. 2013; Johnston et al. 2013; Coccato et al. 2015; Katkov et al. 2016; Morelli et al. 2017; Mitzkus et al. 2017).
We note that, in this work, we also adopted a definition of CRDs based on the circularity parameter, similarly to Abadi et al. (2003), which is more restrictive with respect to the definition of CRDs adopted by Khoperskov et al. 2021 based on the circular velocity. For this reason, in this work, we find that CRDs in late-type galaxies generally comprise a small fraction of the stellar mass with respect to the whole disk, which is more consistent with the results of Jagvaral et al. 2022 based on the TNG100 simulations.